OpenStudy (anonymous):

x ^{2} +y ^{2} -2x +4y -4 =0 how do I find the center and radius of this equation?

OpenStudy (phi):

Do you know how to complete the square?

OpenStudy (phi):

If so, you would complete the square for x^2-2x and do the same for y^2+4y you want to get the equation in the form (x-a)^2 + (y-b)^2 = r^2 (a,b) will be the center, and r the radius

OpenStudy (phi):

@myko have you read the Code of Conduct?

OpenStudy (anonymous):

what did myko do?

OpenStudy (anonymous):

he thinks I willanswer your question

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

I just completed the square, and I got the correct center points but the wrong radius. I am not sure what I did wrong though..

OpenStudy (anonymous):

show your work .

OpenStudy (anonymous):

write it like this: x^2-2x+1+y^2+4y+4-4-4-1=0

OpenStudy (anonymous):

by the way @phi, maybe you can sugest me any good webpage on lebesgue mesure?

OpenStudy (phi):

@myko I'm engineer not a mathematician

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

\[(x ^{2}-2x) + (y ^{2} +4y) =4\] so then I divided -2 by 2, and then squared it which gave me 1 and then I divided 4 by 2 and squared it which then gave me 4. so now the equation looks like \[(x ^{2}-2x-1) + (y ^{2} +4y +2)=5\] I then factored which gave me \[(x-1) + (y-(-2) =5\] so the center would be (1,-2) and the radius would be 5 but my book says that the radius is 3...

OpenStudy (anonymous):

circle equation has a form : \[(x-a)^2+(x-b)^2=c^2\] where (a,b) would be the center and c the radius

OpenStudy (anonymous):

that's why you need to compleate the squeres

OpenStudy (anonymous):

you expression in parenthesis is not a square

OpenStudy (phi):

remember when you square -1 you get +1 you want a +1 not a -1 in the x term

OpenStudy (anonymous):

what do I do then?

OpenStudy (anonymous):

x^2-2x +a = (x - sqrt(a))^2

OpenStudy (anonymous):

what should be a=?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

right, now do the same for y^2+4y +b=(y - sqrt(b))^2

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

sry should be (x + sqrt(a))^2 but 4 is right

OpenStudy (anonymous):

now remmember to add/substract what you added to compleat the square so the expression doesn't change

OpenStudy (anonymous):

so now i have (x+1) + (x+2) = 9

OpenStudy (phi):

don't forget the exponents (the 2's)

OpenStudy (anonymous):

sorry i meant (x+1) + (y+2) =9

OpenStudy (anonymous):

You should have something like this: (x^2-2x+1)+(y^2+4y+4)-4-4-1=0

OpenStudy (anonymous):

which equals to: what you wrote, just with squares: (x-1)^2 + (y+2)^2 =9

OpenStudy (anonymous):

i have (x^2 - 2x +1) + (y^2 +4y +4), but not -4-4-1.. where did that come from?

OpenStudy (anonymous):

never mind I see where it came from

OpenStudy (anonymous):

if you add 1 to compleat the first square you also have to substract it, so the equation stays the same. Same for the 4 in the 2º square

OpenStudy (anonymous):

so your radius would be?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but now my center is wrong.. i have (1,2) when it is supposed to be (1,-2)

OpenStudy (anonymous):

it's right: (x^2-2x+1)+(y^2+4y+4)-4-4-1=0 equals to (x-1)^2 + (y+2)^2 =9

OpenStudy (anonymous):

so center (1,-2)

OpenStudy (anonymous):

why is 2 negative?

OpenStudy (anonymous):

remmeber what i wrote about how circle equation looks like, many comments ago...

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

so, that's why

OpenStudy (anonymous):

(y+2)^2 is same as (y-(-2))^2

OpenStudy (anonymous):

Oh okay. Thank you.

OpenStudy (anonymous):

yw